OpenStudy (anonymous):

How many solutions exist for the following system of equations: 4x + y = 7 and 3x - y = 0 a. no solution b. one solution c. infinitely many d. cannot be determined

OpenStudy (anonymous):

Someone please help, I have extra credit to do by 12 and I really need the points for my grade or I will fail the course

OpenStudy (anonymous):

you can calculate 3x=y 4x+3x=7 x=1 y=3*1=3

OpenStudy (anonymous):

Which of those should i calculate?